OpenStudy (anonymous):

solve by the substitution method 3x+3y=-6 -4x+y=28

OpenStudy (anonymous):

I don't understand it

OpenStudy (anonymous):

have you tryed using the matrix method it works the same way as substitution but its much easier

OpenStudy (anonymous):

I don't know how to do that method...

OpenStudy (anonymous):

if you have a scientific calculator then you could do it

OpenStudy (whpalmer4):

Substitution is easy enough, and good practice for when you don't have something else to solve it for you! \[3x + 3y = -6\]\[-4x + y = 28\] What could you do to one of those equations to express one of the variables in terms of the other? Looks like the second one is easy to solve for y, so let's do that: \[y=4x+28\] (just added 4x to each side) Now substitute (4x+28) wherever you see y in the other equation, which will give you an equation only in terms of x: \[3x+3y= 3x + 3(4x+28) = 15x + 84 = -6\] Subtract 84 from both sides and we get \[15x = -90\]so\[x=-6\] Now use x=-6 in the second equation to solve for y. \[-4(-6) + y = 28\] y = ?

OpenStudy (anonymous):

y=1.4?

OpenStudy (whpalmer4):

\[-4(-6) + y = 28\] \[24 + y = 28\] \[y = ?\]